Derive E=v/d class 12 CBSE
Plz help can anyine solve this !

Grade:12

1 Answers

Eshan
5 years ago
Dear student,

This relation is valid only when electric field through the distance d is constant.

From the definition of potential,\Delta V=-\int Edx
Hence\Delta V=-\int_0^d Edx
For constant E,\Delta V=-E\int_0^d dx=-Ed
HenceE=-\dfrac{V}{d}